# Denoising Refinement Diffusion Models for Simultaneous Generation of Multi-scale Mobile Network Traffic

### v1

@@ III-C Canonical Diffusion Models
The core idea of canonical diffusion models [54, 55, 56] is to construct a forward diffusion process that gradually perturbs the original data into Gaussian noise, while simultaneously training a learnable reverse denoising process to recover the original data. According to Ho et al. [54], the forward diffusion process can be obtained as:
@@
In the reverse process, the model denoises from a standard Gaussian noise $\boldsymbol{x}^{(N-1)}=\boldsymbol{\epsilon}$ and performs iterative reverse sampling according to [54]:
@@ IV-B1 Diffusion Noise Prior Guidance
In Section III-C, Eq. 3 presents the general form of the forward process in diffusion models. Inspired by [57], suppose the real data $\boldsymbol{x}^{(0)}$ contains a fixed, prior-known component $\boldsymbol{H}$, such that $\boldsymbol{x}^{(0)}=\boldsymbol{H}+\Delta\boldsymbol{x}^{(0)}$, where $\Delta\boldsymbol{x}^{(0)}$ denotes the residual between $\boldsymbol{x}^{(0)}$ and $\boldsymbol{H}$. Then, Eq. 3 can be rewritten as:
@@ IV-C1 Noise intensity schedules
In DRDM, we need to constrain the intermediate states at different stage nodes of the denoising process to represent network traffic with specific spatiotemporal scales. The conventional CN causes all diffusion steps, except for the initial noise-adding point, to contain a certain level of random noise, leading to a deviation between the intermediate states at these nodes and the distributions of traffic at corresponding scales. To address this issue, we propose SN. It linearly schedules $\hat{\alpha}$ within each stage while maintaining overall continuity in $(0,1)$:
@@ IV-C2 Noise adding schedules
To effectively constrain the intermediate states during the denoising process, we explicitly provide samples at the stage nodes of the noise-adding process. As shown in Fig. 7, the SA strategy adopts segmented noise addition. Suppose our target is to generate network traffic at four spatiotemporal scales $s_{1}t_{1},~s_{2}t_{2},~s_{3}t_{3},~s_{4}t_{4}$, ranging from coarse to fine, where $s_{4}$ represents the smallest spatial scale and $t_{4}$ denotes the finest temporal resolution. In HNaP, the prior known multi-level network traffic $\boldsymbol{x}_{s_{4}t_{4}},~\boldsymbol{x}_{s_{3}t_{3}},~\boldsymbol{x}_{s_{2}t_{2}},~\boldsymbol{x}_{s_{1}t_{1}}$ are respectively used as the noise-adding starting points for each stage. That is, in Eq. (11), $\boldsymbol{x}^{(0)}_{s_{\delta_{k}}t_{\tau_{k}}}\neq\boldsymbol{x}^{(0)},~k\leq K-1$,
@@ V Evaluation
To evaluate the performance of ZoomDiff on generating multi-scale mobile network traffic with various resolutions, we conduct experiments on a real-world mobile network traffic dataset collected from Nanning, China, and compare it against 10 state-of-the-art baselines.
@@ V-A1 Dataset
We use a real-world mobile network traffic dataset from Nanning, China, which includes mobile traffic records over one week at four layers: BS, Cell, 100m grid, and 50m grid. Due to the limited accessibility of multi-level traffic data, traffic at the BS, Cell, and 100m grid scales is aggregated from the 50m grid traffic according to their respective coverage areas. The finest temporal resolution is 1 hour. Divided by the Yongjiang River, the dataset is split into a South Bank sub-dataset and a North Bank sub-dataset. The South Bank sub-dataset contains 1.4k BSs, 5.7k cells, 23k 100m grids, and 92k 50m grids, while the North Bank subset includes 2.8k BSs, 11k cells, 45k 100m grids, and 182k 50m grids. The dataset is provided by China Mobile.
@@ V-A3 Baselines
We select a total of 10 state-of-the-art baseline models across four categories: i) Basic forecasting models: XGBoost [58] and Transformer [59]; ii) Domain-specific models for network traffic generation: KSTDiff [38], NetDiffus [39], 5GT-GAN [33], and keGAN [35]; iii) Classic diffusion models: CSDI [55] and DiT [56]; iv) Time series generation models: TimeGAN [60], and Diffusion-TS [61].
@@ V-C Multi-resolution Generation
Fig. 10 provides a performance comparison of intermediate denoising states across different resolution hierarchy settings for CSDI, DiT, and ZoomDiff on the South Bank, Nanning dataset. To intuitively illustrate the designed RRDP for DRDM, Fig. 11 presents the temporal waveforms of intermediate denoising states of CSDI and ZoomDiff (S4T4) extracted at $n=450$, 300300, $150,$ and 0. It can be observed that at $n=0$ (completely denoised), the performances of CSDI, DiT, and ZoomDiff are comparable. However, the intermediate denoising states of CSDI and DiT appear relatively chaotic. In contrast, ZoomDiff is able to produce accurate generation results at different stages of the same denoising process. These intermediate states correspond to network traffic with varying spatio-temporal resolutions, rather than noisy and disordered data.
These findings demonstrate that ZoomDiff can effectively regularize the intermediate states of the denoising process without compromising the final high-resolution generation accuracy, which enables the direct extraction of coarser-resolution network traffic from intermediate states. This design innovatively overcomes the limitation of canonical diffusion models, which can only generate data at a single resolution, and provides a unified paradigm for multi-resolution mobile network traffic generation.
@@ V-D1 Diffusion Noise Prior Guidance
As illustrated in Section IV-B2 and Section IV-B3, we employ a prior noise guidance to allow low-resolution network traffic to guide the generation of high-resolution traffic. To achieve Eq. 11, we use a Multi-Layer Perceptron (MLP) to map the predicted residual noise $\boldsymbol{\epsilon}$ and prior noise $\mathcal{P}_{s_{\delta_{k-1}}t_{\tau_{k-1}}}$ in the $D$-dimensional latent space for concatenation. Fig. 12 shows the impact on high-resolution traffic generation accuracy when removing the prior noise guidance or varying the latent noise dimension $D$. It can be observed that the prior noise guidance is essential for DRDM, as the limited number of diffusion steps in one stage is insufficient to fully learn the complex spatio-temporal patterns of high-resolution traffic. Increasing the latent dimension of the noise helps the model better disentangle and fuse the spatio-temporal patterns embedded in the random and prior noise. Experiments show that the performance converges when $D$ reaches 32, indicating that further increasing the dimension provides no additional information gain in the noise fusion process.
@@ V-D2 Spatiotemporal Positional Encoding
In this section, we investigate the role of TPE and SPE in helping the model understand the temporal and spatial characteristics of network traffic. We conduct an ablation study targeting this encoding, with the results shown in Fig. 13. The outcome aligns with our expectations: the introduction of both TPE and SPE provides positive gains to the model.
@@ V-F Efficiency Analysis
In large-scale mobile data management, efficient modeling and generation of spatiotemporal data are crucial. ZoomDiff enables joint generation of multi-scale mobile traffic, producing network traffic at multiple network layers with different spatial scales and temporal resolutions through a single denoising process of one diffusion model. The time complexity of diffusion models mainly arises from the multi-step denoising process; while traditional diffusion models can only generate data at a single spatiotemporal resolution, ZoomDiff improves the utilization efficiency of the denoising inference process.
To verify the efficiency advantage of ZoomDiff, we conduct a four-level network traffic generation task (BS, cell, 100m grid, 50m grid) for an area containing 500 BSs, including both pre-training and inference. For simplicity, the temporal resolution of traffic at each level is fixed at 1 hour. The batch size is set to 4, and a single NVIDIA A100 GPU is used. Table V presents the comparison results between ZoomDiff and representative diffusion models (CSDI, DIT) in terms of parameter scale, training speed, and inference speed. It can be observed that ZoomDiff, while maintaining a small number of parameters (0.504M), achieves remarkable acceleration performance. Its training time is reduced from 22.9 s/epoch in CSDI to 12.25 s/epoch, representing a 46.5% improvement in training efficiency; and its inference time is reduced from 896.6 s/epoch in CSDI to 446.2 s/epoch, representing a 50.2% improvement in inference efficiency. These results demonstrate the superior efficiency of ZoomDiff when generating multi-scale mobile traffic.
@@
VI Generative Management of Multi-scale Mobile Traffic Data
With the increasing demand for joint optimization across multi-level mobile networks, both operators and researchers urgently require optimization training based on massive mobile data. However, due to the privacy of traffic data and the incomplete deployment of communication infrastructure, multi-level network traffic data, especially fine-grained mobile traffic, remains difficult to obtain in many regions. Therefore, implementing full-lifecycle management of multi-scale mobile traffic based on generative AI has become an urgent need. Designed for joint generation of multi-resolution traffic, ZoomDiff demonstrates unique advantages during mobile data management, including high accessibility of input conditions, high accuracy, and high efficiency in training and inference. This section analyzes the performance of ZoomDiff from the perspectives of two key operations in mobile traffic data management: Traffic Refinement and cross-city generalization.
VI-A Management of Traffic Refinement
The refinement of coarse-grained traffic is an important task in mobile data management. User-level network planning often requires the estimation of grid-level network loads. However, operators can typically only obtain large-scale traffic data such as BS-level or cell-level traffic, while finer-grained traffic estimation relies on point sampling and drive testing, which suffer from limited accuracy and high cost. The pre-trained ZoomDiff model is capable of understanding the correlations among multi-scale traffic. By setting the denoising starting point of the reverse process to specific coarse-scale data, ZoomDiff can directly refine BS or cell traffic into grid-level fine-grained traffic. Table VI presents the results of refining BS traffic into 50m grid traffic using different methods. In Table VI, ZoomDiff (zero-shot) directly inputs BS-level traffic into the corresponding denoising intermediate state during the reverse process, while CSDI and ZoomDiff were retrained for the traffic refinement task. The results show that ZoomDiff (zero-shot) achieves accuracy comparable to the retrained models, demonstrating its excellent performance in the management of traffic refinement.
@@ VI-B Generalization Validation
Mobile data from different cities and operators is often difficult to share due to privacy concerns, which poses obstacles to multi-city mobile data management that is required for the testing of optimization strategies. Utilizing generative models to generalize the spatiotemporal distribution patterns of multi-scale traffic from known cities to unknown ones can effectively address this problem. To validate the cross-city and multi-scenario generalization capability of ZoomDiff, we conducted transfer performance tests on two cities, Nanning and Shanghai, using traffic datasets from two levels of base stations: Macro-BS (MBS) and Pico-BS (PBS). The generation results of 1-hour granularity traffic at the PBS level are shown in Table VII. It can be observed that, compared with CSDI, ZoomDiff exhibits a relative accuracy advantage in both single-city training and cross-city transfer scenarios.
@@ VII Conclusion
In this paper, we propose ZoomDiff, a diffusion-based multi-scale mobile traffic generation model. ZoomDiff employs customized multi-resolution diffusion models, DRDM. It achieves joint modeling of multi-layer network traffic with different spatial and temporal resolutions through multi-stage noise-adding and denoising processes. By leveraging a prior noise guidance mechanism and multi-stage denoising, ZoomDiff can simultaneously output multi-scale mobile traffic within a single denoising process by controlling the intermediate denoising states. We further explored effective noise scheduling strategies and the resolution greediness preference for DRDM. Experiments on real-world datasets demonstrate that ZoomDiff exhibits excellent performance in generating multi-scale network traffic. In addition, ZoomDiff effectively enhances the management of traffic refinement and shows strong cross-city generalization capability. This provides an effective solution for efficient generative management of multi-scale mobile traffic data.
AI-Generated Content Acknowledgement
No generative AI tools were used in the conception, writing, or experimentation of this paper. All text, figures, images, and code in the paper and related materials were created entirely by the authors without the assistance of generative AI tools.

### v2

@@ III-C Canonical Diffusion Models
The core idea of canonical diffusion models [DDPM, CSDI, DiT] is to construct a forward diffusion process that gradually perturbs the original data into Gaussian noise, while simultaneously training a learnable reverse denoising process to recover the original data. According to Ho et al. [DDPM], the forward diffusion process can be obtained as:
@@
In the reverse process, the model denoises from a standard Gaussian noise $\boldsymbol{x}^{(N-1)}=\boldsymbol{\epsilon}$ and performs iterative reverse sampling according to [DDPM]:
@@ IV-B1 Diffusion Noise Prior Guidance
In Section III-C, Eq. 3 presents the general form of the forward process in diffusion models. Inspired by [2025arXiv250113794S], suppose the real data $\boldsymbol{x}^{(0)}$ contains a fixed, prior-known component $\boldsymbol{H}$, such that $\boldsymbol{x}^{(0)}=\boldsymbol{H}+\Delta\boldsymbol{x}^{(0)}$, where $\Delta\boldsymbol{x}^{(0)}$ denotes the residual between $\boldsymbol{x}^{(0)}$ and $\boldsymbol{H}$. Then, Eq. 3 can be rewritten as:
@@ IV-C1 Noise intensity schedules
In DRDM, we need to constrain the intermediate states at different stage nodes of the denoising process to represent mobile traffic with specific spatiotemporal scales. The conventional CN causes all diffusion steps, except for the initial noise-adding point, to contain a certain level of random noise, leading to a deviation between the intermediate states at these nodes and the distributions of traffic at corresponding scales. To address this issue, we propose SN. It linearly schedules $\hat{\alpha}$ within each stage while maintaining overall continuity in $(0,1)$:
@@ IV-C2 Noise adding schedules
To effectively constrain the intermediate states during the denoising process, we explicitly provide samples at the stage nodes of the noise-adding process. As shown in Fig. 7, the SA strategy adopts segmented noise addition. Suppose our target is to generate mobile traffic at four spatiotemporal scales $s_{1}t_{1},~s_{2}t_{2},~s_{3}t_{3},~s_{4}t_{4}$, ranging from coarse to fine, where $s_{4}$ represents the smallest spatial scale and $t_{4}$ denotes the finest temporal resolution. In HNaP, the prior known multi-level mobile traffic $\boldsymbol{x}_{s_{4}t_{4}},~\boldsymbol{x}_{s_{3}t_{3}},~\boldsymbol{x}_{s_{2}t_{2}},~\boldsymbol{x}_{s_{1}t_{1}}$ are respectively used as the noise-adding starting points for each stage. That is, in Eq. (11), $\boldsymbol{x}^{(0)}_{s_{\delta_{k}}t_{\tau_{k}}}\neq\boldsymbol{x}^{(0)},~k\leq K-1$,
@@ V Evaluation
To evaluate the performance of ZoomDiff on generating multi-scale mobile traffic with various resolutions, we conduct experiments on a real-world mobile traffic dataset collected from Nanning, China, and compare it against 10 state-of-the-art baselines.
@@ V-A1 Dataset
We use a real-world mobile traffic dataset from Nanning, China, which includes mobile traffic records over one week at four layers: BS, Cell, 100m grid, and 50m grid. Due to the limited accessibility of multi-level traffic data, traffic at the BS, Cell, and 100m grid scales is aggregated from the 50m grid traffic according to their respective coverage areas. The finest temporal resolution is 1 hour. Divided by the Yongjiang River, the dataset is split into a South Bank sub-dataset and a North Bank sub-dataset. The South Bank sub-dataset contains 1.4k BSs, 5.7k cells, 23k 100m grids, and 92k 50m grids, while the North Bank subset includes 2.8k BSs, 11k cells, 45k 100m grids, and 182k 50m grids. The dataset is provided by China Mobile.
@@ V-A2 Evaluation Metrics
MAE and RMSE assess the overall accuracy of spatiotemporal sequence generation, while spRMSE and PSNR evaluate the accuracy of spatial distribution.
@@ V-A3 Baselines
We select a total of 10 state-of-the-art baseline models across four categories: i) Basic forecasting models: XGBoost [XGBoost] and Transformer [Transformer]; ii) Domain-specific models for mobile traffic generation: KSTDiff [KSTDiff], NetDiffus [NetDiffus], 5GT-GAN [5GT-GAN], and keGAN [keGAN]; iii) Classic diffusion models: CSDI [CSDI] and DiT [DiT]; iv) Time series generation models: TimeGAN [TimeGAN], and Diffusion-TS [Diffusion-TS].
@@ V-C Multi-resolution Generation
Fig. 10 provides a performance comparison of intermediate denoising states across different resolution hierarchy settings for CSDI, DiT, and ZoomDiff on the South Bank, Nanning dataset. To intuitively illustrate the designed RRDP for DRDM, Fig. 11 presents the temporal waveforms of intermediate denoising states of CSDI and ZoomDiff (S4T4) extracted at $n=450$, 300300, $150,$ and 0. It can be observed that at $n=0$ (completely denoised), the performances of CSDI, DiT, and ZoomDiff are comparable. However, the intermediate denoising states of CSDI and DiT appear relatively chaotic. In contrast, ZoomDiff is able to produce accurate generation results at different stages of the same denoising process. These intermediate states correspond to mobile traffic with varying spatio-temporal resolutions, rather than noisy and disordered data.
These findings demonstrate that ZoomDiff can effectively regularize the intermediate states of the denoising process without compromising the final high-resolution generation accuracy, which enables the direct extraction of coarser-resolution mobile traffic from intermediate states. This design innovatively overcomes the limitation of canonical diffusion models, which can only generate data at a single resolution, and provides a unified paradigm for multi-resolution mobile traffic generation.
@@ V-D1 Diffusion Noise Prior Guidance
As illustrated in Section IV-B2 and Section IV-B3, we employ a prior noise guidance to allow low-resolution mobile traffic to guide the generation of high-resolution traffic. To achieve Eq. 11, we use a Multi-Layer Perceptron (MLP) to map the predicted residual noise $\boldsymbol{\epsilon}$ and prior noise $\mathcal{P}_{s_{\delta_{k-1}}t_{\tau_{k-1}}}$ in the $D$-dimensional latent space for concatenation. Fig. 12 shows the impact on high-resolution traffic generation accuracy when removing the prior noise guidance or varying the latent noise dimension $D$. It can be observed that the prior noise guidance is essential for DRDM, as the limited number of diffusion steps in one stage is insufficient to fully learn the complex spatio-temporal patterns of high-resolution traffic. Increasing the latent dimension of the noise helps the model better disentangle and fuse the spatio-temporal patterns embedded in the random and prior noise. Experiments show that the performance converges when $D$ reaches 32, indicating that further increasing the dimension provides no additional information gain in the noise fusion process.
@@ V-D2 Spatiotemporal Positional Encoding
In this section, we investigate the role of TPE and SPE in helping the model understand the temporal and spatial characteristics of mobile traffic. We conduct an ablation study targeting this encoding, with the results shown in Fig. 13. The outcome aligns with our expectations: the introduction of both TPE and SPE provides positive gains to the model.
@@ V-F Efficiency Analysis
Efficient modeling and generation of spatiotemporal data are crucial for large-scale mobile network applications. ZoomDiff achieves joint generation of multi-scale mobile traffic by producing mobile traffic at multiple layers with different spatial scales and temporal resolutions through a single denoising process of one diffusion model. The time complexity of diffusion models mainly arises from their iterative denoising steps. Traditional diffusion models can only generate data at a single spatiotemporal resolution, leading to redundant computations when multi-scale data are required. In contrast, ZoomDiff enhances the utilization efficiency of the denoising process by enabling multiple resolutions to be generated simultaneously within a single inference cycle, thereby significantly improving computational efficiency.
To verify the efficiency advantage of ZoomDiff, we conduct a four-level mobile traffic generation task (BS, cell, 100m grid, 50m grid) for an area containing 500 BSs, including both pre-training and inference. For simplicity, the temporal resolution of traffic at each level is fixed at 1 hour. The batch size is set to 4, and a single NVIDIA A100 GPU is used. Table V presents the comparison results between ZoomDiff and representative diffusion models (CSDI, DIT) in terms of parameter scale, training speed, and inference speed. It can be observed that ZoomDiff, while maintaining a small number of parameters (0.504M), achieves remarkable acceleration performance. Its training time is reduced from 22.9 s/epoch in CSDI to 12.25 s/epoch, representing a 46.5% improvement in training efficiency; and its inference time is reduced from 896.6 s/epoch in CSDI to 446.2 s/epoch, representing a 50.2% improvement in inference efficiency. These results demonstrate the superior efficiency of ZoomDiff when generating multi-scale mobile traffic.
@@
VI Case Study
In this section, we discuss two key aspects of ZoomDiff: its capability for traffic refinement from coarse (BS level) to fine granularity, and its generalization ability across different cities and network environments.
VI-A Refinement of BS-level Traffic Data
The refinement of coarse-grained traffic into finer resolutions is a crucial task for enhancing the interpretability and usability of mobile network data. In real-world scenarios, operators typically obtain only large-scale traffic data such as BS-level or cell-level measurements, while finer-grained traffic at the grid level often depends on point sampling or drive testing, which are costly and prone to errors. Leveraging its understanding of multi-scale correlations, the pre-trained ZoomDiff model can refine BS or cell-level traffic into grid-level traffic through a zero-shot inference process. Specifically, by initializing the reverse diffusion process with coarse-scale traffic as the denoising starting point, ZoomDiff can progressively generate finer-grained spatial distributions without additional retraining. Table VI presents the comparison results for refining BS traffic into 50m-grid traffic using different approaches. The results show that ZoomDiff (zero-shot) achieves accuracy close to that of retrained models, significantly outperforming traditional baselines such as CSDI. This demonstrates that ZoomDiff effectively captures the hierarchical correlations in mobile traffic and supports efficient fine-grained refinement.
@@ VI-B Generalization Validation
Mobile traffic patterns often differ significantly across cities due to variations in population density, infrastructure deployment, and user behavior. However, it is impractical to collect sufficient training data for every city due to privacy and deployment constraints. Therefore, the ability to generalize learned traffic generation models across cities is essential. To evaluate the cross-city generalization of ZoomDiff, we conducted transfer experiments between Nanning and Shanghai using Macro-BS (MBS) and Pico-BS (PBS) datasets. In the transfer setting, models trained on Nanning data were tested on Shanghai data without fine-tuning. The results, shown in Table VII, indicate that ZoomDiff achieves consistently better performance than CSDI in both single-city training and cross-city transfer. Notably, the degradation in accuracy during transfer is minor, demonstrating that ZoomDiff effectively captures universal spatiotemporal dependencies that are transferable across diverse urban environments.
@@ VII Conclusion
In this paper, we propose ZoomDiff, a diffusion-based model for multi-scale mobile traffic generation. ZoomDiff employs customized multi-resolution diffusion models, DRDM. It achieves joint modeling of multi-layer mobile traffic with different spatial and temporal resolutions through multi-stage noise-adding and denoising processes. By leveraging a prior noise guidance mechanism and hierarchical denoising stages, ZoomDiff can generate multi-scale mobile traffic within a single denoising process by controlling intermediate states. Furthermore, we investigate effective noise scheduling strategies and analyze the resolution greediness preference in DRDM, which enhances the stability and scalability of the model. Extensive experiments on real-world datasets demonstrate that ZoomDiff achieves at least an 18.4% improvement over state-of-the-art baselines in generating multi-scale mobile traffic and shows strong cross-city generalization capability. These results highlight ZoomDiff as a promising foundation for future research in generative modeling of large-scale mobile network dynamics.